# Pseudo-Simulation for Autonomous Driving

### v2

@@ 3.1 Stage 1: Initial Observations
BEV Simulation. The 4-second trajectory predicted by the agent is executed using a kinematic bicycle model [41] and an LQR controller [42] at 10Hz. The trajectory is committed for the entire simulation horizon, and no closed-loop feedback is provided to the agent during this time. Unlike related prior work [8], which uses non-reactive traffic to simplify implementation (i.e., neighboring vehicles follow their recorded trajectories without reacting to the ego agent), we improve the simulation realism with reactive traffic. Background vehicles (represented as oriented bounding boxes) respond to the ego agent using a rule-based planner called the Intelligent Driver Model (IDM) [43].
@@ 4.1 How well-aligned is pseudo-simulation with closed-loop evaluation?
Figure 3: Correlations. (a) Correlation between the default pseudo-simulation metric (EPDMS) and the closed-loop score (CLS) for a set of 37 rule-based and 46 learned planners. We further compare (b) single ($1\mathsf{x}$) vs. two stage ($2\mathsf{x}$) evaluation, (c) Gaussian weight variances, (d) Stage 1 and 2 aggregation methods, and (e) synthetic observation densities. Defaults in bold-underline.
@@
In Fig. 3 (b), we compare single-stage open-loop simulation (at 4 and 8 seconds) to our two-stage pseudo-simulation variant. The two-stage setup achieves significantly higher alignment, reaching a Pearson correlation of $r=0.89$ (corresponding to $R^{2}=0.8$), compared to $r=0.83$ ($R^{2}=0.7$) for the single-stage baselines. Furthermore, compared to standard reactive closed-loop evaluation, our pseudo-simulation method exposes a wider range of potential failures. This typically results in lower average EPDMS values compared to CLS values. By injecting synthetic deviations, pseudo-simulation effectively reveals edge cases that might not be encountered during standard testing.
Within Stage 2, we assess the impact of the weighting used to combine scores across synthetic viewpoints.  Fig. 3 (c) shows the correlation with closed-loop scores for different kernel variances. We observe that smaller variances lead to improved results. $\sigma^{2}=0.05$ and our default configuration of $\sigma^{2}=0.1$ give the highest correlations. In additional experiments (included in the supplementary material), we find that other approaches, such as simple averaging, $k$-nearest neighbors ($k$-NN), and hybrid $k$-NN/Gaussian weighting are less effective than our default configuration.
To combine metrics across stages, we compare multiplicative aggregation, aggregation by the arithmetic mean, and a hybrid strategy where penalty metrics (e.g., collision and drivable area compliance) are multiplied while the remaining terms are averaged.  Fig. 3 (d) summarizes the correlation of each strategy with closed-loop scores. Multiplicative aggregation shows both higher linear and rank correlation than the other approaches. This outcome is likely because most subscores are binary (i.e., 0 or 1). Consequently, multiplication appears to be a more suitable method for estimating the overall score for an 8-second interval based on two 4-second segments.
@@ 4.2 What new challenges and insights does our leaderboard provide?
Benchmark. Our public NAVSIM v2 leaderboard features challenging driving scenarios (e.g. unprotected turns and dense traffic, see Fig. 2). It uses a subset of nuPlan that we refer to as navhard, involving 450 Stage 1 and 5462 Stage 2 observations. Further details are in the supplementary.
Metric Stage CV [8] MLP [8] LTF [11] PDM-C [12]
NC $\uparrow$ S1 88.8 93.2 96.2 94.4
S2 83.2 77.2 77.7 88.1
DAC $\uparrow$ S1 42.8 55.7 79.5 98.8
S2 59.1 51.9 70.2 90.6
DDC $\uparrow$ S1 70.6 86.6 99.1 100
S2 76.5 74.4 84.2 96.3
TLC $\uparrow$ S1 99.3 99.3 99.5 99.5
S2 98.0 98.2 98.0 98.5
EP $\uparrow$ S1 77.5 81.2 84.1 100
S2 71.3 77.1 85.1 100
TTC $\uparrow$ S1 87.3 92.2 95.1 93.5
S2 81.1 75.0 75.6 83.1
LK $\uparrow$ S1 78.6 83.5 94.2 99.3
S2 47.9 40.8 45.4 73.7
HC $\uparrow$ S1 97.1 97.5 97.5 87.7
S2 97.1 97.8 95.7 91.5
EC $\uparrow$ S1 60.4 77.7 79.1 36.0
S2 61.9 79.8 75.9 25.4
EPDMS $\uparrow$ 10.9 12.7 23.1 51.3
Table 2: navhard leaderboard.
We select four baseline planners with varying input modalities. These include the Constant Velocity (CV) and Ego-history MLP from [8], Latent TransFuser (LTF) [11], the strongest public image-only planner for nuPlan, and, PDM-Closed (PDM-C) [12], the best privileged planner on nuPlan.
Results. Table 2 presents the detailed subscores for each planner, broken down by Stage 1 (original observations) and Stage 2 (synthetic observations). In terms of overall performance, PDM-Closed achieves the highest combined EPDMS of 51.3, followed by LTF (23.1), MLP (12.7), and CV (10.9). Comparing performance across stages, we observe a general drop in subscores from Stage 1 to Stage 2 for all methods (particularly for LTF), suggesting their sensitivity to the distribution shifts introduced in Stage 2. Notably, while PDM-Closed excels in most metrics, it exhibits lower performance in comfort metrics such as HC and EC. Our evaluation on navhard reveals this specific failure mode of PDM-Closed, highlighting a trade-off that was overlooked in prior benchmarks. We host navhard as a public leaderboard.
@@ 4.3 Does the proposed neural rendering yield sufficient visual fidelity?
Benchmark. To assess the fidelity of our synthetic observations, we evaluate the impact of our neural rendering on the downstream perception and planning performance of a pre-trained model using the navhard dataset. Specifically, we employ the LTF model from Table 2. Although primarily an end-to-end planner, LTF outputs intermediate Bird’s Eye View (BEV) segmentations and, crucially, was trained only on real-world data. This allows us to measure the domain gap introduced by our rendering: we evaluate LTF’s performance on synthetic data and compare it to its performance on real data. We use mean Intersection over Union (mIoU) over the drivable area, walkway, and vehicle classes output by LTF to evaluate BEV perception, and the EPDMS metric to evaluate planning.
@@ 5 Conclusion
Human-flag Filtering. Our filtering strategy disregards rule violations also committed by human experts. While this helps reduce false positives, it could also risk overlooking important failure and edge cases, since human driving is not always a gold standard for safety. Future work could further refine the human-flag filtering and explore this trade-off to ensure more reliable evaluation.
@@ Acknowledgments
This work was supported by the ERC Starting Grant LEGO-3D (850533), the DFG EXC number 2064/1 - project number 390727645, the German Federal Ministry of Education and Research: Tübingen AI Center, FKZ: 01IS18039A and the German Federal Ministry for Economic Affairs and Climate Action within the project NXT GEN AI METHODS. We thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Daniel Dauner and Kashyap Chitta. We also thank HuggingFace for hosting our evaluation servers, the team members of OpenDriveLab for their organizational support, Maxim Dolgov for helpful discussions, as well as Napat Karnchanachari and the team from Motional for open-sourcing their dataset and providing us the private test split used in the 2025 NAVSIM Challenge.

### v3

@@ 3.1 Stage 1: Initial Observations
BEV Simulation. The 4-second trajectory predicted by the agent is executed using a kinematic bicycle model [41] and an LQR controller [42] at 10Hz. The trajectory is committed for the entire simulation horizon, and no closed-loop feedback is provided to the agent during this time. Unlike related prior work [8], which uses non-reactive traffic to simplify implementation (i.e., neighboring vehicles follow their recorded trajectories without reacting to the ego agent), we improve the simulation realism with reactive traffic. Background vehicles (represented as oriented bounding boxes) respond to the ego agent using a rule-based planner called the Intelligent Driver Model (IDM) [43]. Pedestrians, static obstacles, and other non-vehicle actors follow their recorded trajectories without reacting to the ego agent.
@@ 4 Results
In this section, we study three key questions to examine our proposed idea of pseudo-simulation.
@@ 4.1 How well-aligned is pseudo-simulation with closed-loop evaluation?
Figure 3: Pseudo-simulation correlates strongly with closed-loop evaluation. Compared to the average displacement error (ADE) and nuPlan open-loop score (OLS) metrics, EPDMS achieves a strong correlation to the closed-loop score (CLS) for a set of 37 rule-based and 46 learned planners.
@@
In Fig. 4 (a), we compare single-stage open-loop simulation (at 4 and 8 seconds) to our two-stage pseudo-simulation variant. The two-stage setup achieves significantly higher alignment, reaching a Pearson correlation of $r=0.89$ (corresponding to $R^{2}=0.8$), compared to $r=0.83$ ($R^{2}=0.7$) for the single-stage baselines. Furthermore, compared to standard reactive closed-loop evaluation, our pseudo-simulation method exposes a wider range of potential failures. This typically results in lower average EPDMS values compared to CLS values. By injecting synthetic deviations, pseudo-simulation effectively reveals edge cases that might not be encountered during standard testing.
Within Stage 2, we assess the impact of the weighting used to combine scores across synthetic viewpoints.  Fig. 4 (b) shows the correlation with closed-loop scores for different kernel variances. We observe that smaller variances lead to improved results. $\sigma^{2}=0.05$ and our default configuration of $\sigma^{2}=0.1$ give the highest correlations. In additional experiments (included in the supplementary material), we find that other approaches, such as simple averaging, $k$-nearest neighbors ($k$-NN), and hybrid $k$-NN/Gaussian weighting are less effective than our default configuration.
Figure 4: EPDMS design choices. We compare the correlation between the pseudo-simulation metric and the closed-loop score (CLS) under several settings. (a) single ($1\mathsf{x}$) vs. two stage ($2\mathsf{x}$) evaluation, (b) Gaussian weight variances, (c) Stage 1 and 2 aggregation methods, and (d) synthetic observation densities. Defaults in bold-underline.
To combine metrics across stages, we compare multiplicative aggregation, aggregation by the arithmetic mean, and a hybrid strategy where penalty metrics (e.g., collision and drivable area compliance) are multiplied while the remaining terms are averaged.  Fig. 4 (c) summarizes the correlation of each strategy with closed-loop scores. Multiplicative aggregation shows both higher linear and rank correlation than the other approaches. This outcome is likely because most subscores are binary (i.e., 0 or 1). Consequently, multiplication appears to be a more suitable method for estimating the overall score for an 8-second interval based on two 4-second segments.
@@ 4.2 What new challenges and insights does our leaderboard provide?
Stage
CV [8]
Ego MLP [8]
LTF [11]
NavFormer
LTFv6 [51]
RAP [52]
ZTRS [53]
GuideFlow [54]
SimScale [55]
DrivoR [56]
PDM-C [12]
NC $\uparrow$ S1 88.8 93.2 96.2 96.2 96.5 97.1 98.8 99.5 99.5 99.1 94.4
S2 83.2 77.2 77.7 85.7 79.8 83.2 91.1 91.4 94.5 92.3 90.5
DAC $\uparrow$ S1 42.8 55.7 79.5 92.4 86.6 94.4 97.5 98.0 99.1 98.2 98.8
S2 59.1 51.9 70.2 81.0 75.5 83.8 90.4 89.5 94.2 91.6 90.6
DDC $\uparrow$ S1 70.6 86.6 99.1 95.7 99.2 98.7 100.0 99.4 99.8 99.3 100
S2 76.5 74.4 84.2 83.5 86.2 87.3 95.7 95.1 95.7 97.3 95.4
TLC $\uparrow$ S1 99.3 99.3 99.5 99.6 99.5 99.7 100.0 99.3 100.0 99.7 99.5
S2 98.0 98.2 98.0 97.6 97.8 98.0 99.0 98.8 99.2 99.0 98.4
EP $\uparrow$ S1 77.5 81.2 84.1 83.8 84.4 83.8 66.6 79.6 69.6 75.4 100
S2 71.3 77.1 85.1 90.1 89.5 86.8 63.5 77.5 75.8 75.7 100
TTC $\uparrow$ S1 87.3 92.2 95.1 96.0 95.1 96.8 98.8 99.3 99.5 98.6 93.5
S2 81.1 75.0 75.6 82.4 76.0 80.3 89.7 89.5 92.8 90.5 86.6
LK $\uparrow$ S1 78.6 83.5 94.2 94.7 94.4 94.6 96.2 94.8 95.7 94.8 99.3
S2 47.9 40.8 45.4 48.2 50.0 52.2 60.4 52.5 60.0 56.0 74.2
HC $\uparrow$ S1 97.1 97.5 97.5 96.4 97.7 96.4 96.6 97.1 95.5 97.5 87.7
S2 97.1 97.8 95.7 94.9 95.2 95.1 97.6 93.5 96.0 98.4 91.9
EC $\uparrow$ S1 60.4 77.7 79.1 60.9 76.4 66.2 44.0 58.2 28.4 70.2 36.0
S2 61.9 79.8 75.9 48.4 66.7 52.4 66.0 51.0 43.2 44.7 29.7
EPDMS $\uparrow$ 11.4 14.1 25.1 34.1 31.9 39.6 48.1 51.5 53.2 54.5 56.6
Table 2: navhard leaderboard. Snapshot from 03/2026.
Benchmark. Our public NAVSIM v2 leaderboard [^1] https://huggingface.co/spaces/AGC2025/e2e-driving-navhard features challenging driving scenarios (e.g. unprotected turns and dense traffic, see Fig. 2). It uses a subset of nuPlan that we refer to as navhard, involving 450 Stage 1 and 5462 Stage 2 observations. The goal of the leaderboard is to ensure standardized evaluation, as subtle differences (e.g., enabling/disabling the human-based filtering mechanism, or using a newer numpy version) can have an effect on the computed metrics. Therefore, we discourage the use of self-reported and unofficial “NAVSIM v2” benchmark splits, such as reporting the EPDMS on the NAVSIM v1 navtest dataset without conducting two-stage pseudo-simulation. Instead, submission to our leaderboard can ensure both consistency and visibility of submissions.
For our analysis, we select five baseline planners with varying input modalities, as well as six community-contributed submissions: LTFv6 [51], RAP [52], ZTRS [53], GuideFlow [54], SimScale [55], and DrivoR [56]. Our baselines include the Constant Velocity (CV) and Ego-history MLP from [8], Latent TransFuser (LTF) [11], the most established baseline for image-only planning for nuPlan, and PDM-Closed (PDM-C) [12], the best privileged planner on nuPlan. To tackle the complex scene types and visual perturbations of navhard, we propose a new simple and robust baseline architecture called NavFormer [^2] https://github.com/OpenDriveLab/NavFormer. This model employs a BEVFormer encoder [57] to process surround-view camera inputs, in contrast to LTF which only encodes forward-facing camera data. Following PARA-Drive [6], object tracking and map segmentation decoders provide auxiliary perception supervision. Finally, NavFormer uses a Hydra-MDP decoder head for planning [58].
Results. Table 2 presents the detailed subscores for each planner, broken down by Stage 1 (original observations) and Stage 2 (synthetic observations). In terms of overall performance, PDM-Closed achieves the highest EPDMS of 56.6, closely followed by recent scalable architectures like DrivoR [56]. Comparing performance across stages, we observe a general drop in subscores from Stage 1 to Stage 2 for all methods, suggesting their sensitivity to the distribution shifts introduced in Stage 2. Notably, while PDM-Closed excels in most metrics, it exhibits lower performance in comfort metrics such as HC and EC. Our evaluation on navhard reveals this specific failure mode of PDM-Closed, highlighting a trade-off that was overlooked in prior benchmarks.
@@ 4.3 Does the proposed neural rendering yield sufficient visual fidelity?
Benchmark. To assess the fidelity of our synthetic observations, we evaluate the impact of our neural rendering on the downstream perception and planning performance of a pre-trained model using the navhard dataset. Specifically, we employ the LTF model from Table 2. Although primarily an end-to-end planner, LTF outputs intermediate Bird’s Eye View (BEV) segmentations and, crucially, is trained only on real-world data. This allows us to measure the domain gap introduced by our rendering: we evaluate LTF’s performance on synthetic data and compare it to its performance on real data. We use mean Intersection over Union (mIoU) over the drivable area, walkway, and vehicle classes output by LTF to evaluate BEV perception, and the EPDMS metric to evaluate planning.
@@ 5 Conclusion
Human Flag Filtering. Our filtering strategy disregards rule violations also committed by human experts. While this helps reduce false positives, it could also risk overlooking important failure and edge cases, since human driving is not always a gold standard for safety. Future work could further refine the human flag filtering and explore this trade-off to ensure more reliable evaluation.
@@ Acknowledgments
This work was supported by the ERC Starting Grant LEGO-3D (850533), the DFG EXC number 2064/1 - project number 390727645, the German Federal Ministry of Education and Research: Tübingen AI Center, FKZ: 01IS18039A and the German Federal Ministry for Economic Affairs and Energy within the project NXT GEN AI METHODS. We thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Daniel Dauner and Kashyap Chitta. We also thank HuggingFace for hosting our evaluation servers, the team members of OpenDriveLab for their organizational support, Maxim Dolgov for helpful discussions, as well as Napat Karnchanachari and the team from Motional for open-sourcing their dataset and providing us the private test split used in the 2025 NAVSIM Challenges.